### v1

# Agent-Oriented Planning in Multi-Agent Systems

Abstract
Through the collaboration of multiple agents possessing diverse expertise and tools, multi-agent systems achieve impressive progress in solving real-world problems. Given the user queries, the meta-agents, serving as the brain within these systems, are required to decompose the queries into multiple sub-tasks that can be allocated to suitable agents capable of solving them, so-called agent-oriented planning. In this study, we identify three critical design principles of agent-oriented planning, including solvability, completeness, and non-redundancy, to ensure that each sub-task is effectively resolved, leading to satisfactory responses to the original queries. These principles further inspire us to propose a novel framework for agent-oriented planning in multi-agent systems, leveraging a fast task decomposition and allocation process followed by an effective and efficient evaluation via a reward model. During the planning process, the meta-agent is also responsible for evaluating the performance of the expert agents, making timely adjustments to the sub-tasks and scheduling as necessary. Besides, we integrate a feedback loop into the proposed framework to further enhance the effectiveness and robustness of such a problem-solving process. Extensive experiments demonstrate the advancement of the proposed framework in solving real-world problems compared to both single-agent systems and existing planning strategies for multi-agent systems.
1 Introduction
In recent years, large language models (LLMs) (Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023) have achieved impressive breakthroughs in natural language understanding and generation, marking a critical advancement in the exploration of artificial general intelligence (AGI). As the capabilities of LLMs progress, LLM-empowered agents (Qin et al., 2023b; Dong et al., 2023; Wang et al., 2024) emerge as key components for integrating expertise and tools to translate these advancements into practical applications effectively. Advancing this paradigm, multi-agent systems (Shen et al., 2024; Chen et al., 2023; Wu et al., 2023b), which involve multiple diverse agents, provide great flexibility and adaptability by leveraging and collaborating on the strengths of various agents, promoting comprehensive solutions to complex real-world problems.
Previous studies (Cai et al., 2023; Qian et al., 2023) propose to define suitable standard operating procedures (SOPs) based on the insights and experiences of human professionals for solving specific tasks, such as software development (Hong et al., 2023) and simulating personality traits (Serapio-García et al., 2023). In these scenarios, multiple agents are assigned to execute tasks following the predefined SOPs, resulting in remarkable successes. However, for multi-agent systems designed to tackle a diverse range of complex real-world problems, there is a critical need for a central entity that can automatically generate task-specific operating procedures and coordinate the activities of various agents (Wang et al., 2024).
This central entity, typically referred to as a meta-agent (a.k.a. a controller or planner), carries two primary responsibilities. Firstly, the meta-agent needs to comprehend user queries and decompose them into several sub-tasks, ensuring that each sub-task can be adequately addressed by a single agent. Secondly, the meta-agent is also expected to assign these sub-tasks to the appropriate agents for execution, enabling the solutions of these sub-tasks to collectively provide comprehensive answers to the original user query.
However, since the meta-agent cannot effectively associate sub-tasks based on the provided agent descriptions, the performance of task decomposition and allocations might be sub-optional. To tackle these challenges, we identify three design principles for an agent-oriented planning framework, including solvability, completeness, and non-redundancy, which further inspire us to propose a novel framework for multi-agent systems.
Specifically, to resolve a user query, the meta-agent first performs fast task decomposition and allocation, which serves as the intermediate results that can be further modified and revised. The proposed framework incorporates a reward model designed to efficiently evaluate the solvability of sub-tasks without requiring actual agent calls. According to the evaluation results, some sub-tasks may be executed by the assigned agents, others may be recognized as inappropriate and be required to perform a replan, while the remaining sub-tasks would be further assessed through the representative works mechanism, which helps determine whether they should be planned in detail or re-described for better aligning the ability of agents. A detector is utilized to identify the missing key information or redundant content in the decomposed sub-tasks, and to provide suggestions to the meta-agent for enhancing the completeness and non-redundancy. Besides, a feedback loop is integrated into the proposed framework to promote ongoing enhancements for the meta-agent.
Extensive experiments are conducted based on a reasoning dataset that needs the collaboration of multiple agents. The comparisons between the proposed agent-oriented planning framework and baseline methods demonstrate the remarkable advancements achieved by the proposed framework. Besides, we conduct an ablation study to confirm the contributions of the different components in the proposed framework, and provide discussions on the potential for further improvements in the performance of agent-oriented planning within multi-agent systems.
2 Preliminaries
Definition of Agent-Oriented Planning
In this study, a multi-agent system involves a series of LLM-empowered agents that possess diverse expertise and tools. These agents are designed to leverage their expertise and tools to resolve specific real-world tasks. For example, a search agent can utilize search engines to retrieve up-to-date information relevant to a given query, while a code agent is capable of generating code and executing the code via a code interpreter. The collaboration among these agents is facilitated by a meta-agent, which serves as the brain of the multi-agent systems. When a user query is submitted to the multi-agent system, the meta-agent sends the relevant requests to the appropriate agents, optimizing for both effectiveness and efficiency. The responses produced by these agents are finally aggregated and synthesized to generate a comprehensive answer to the original user query.
Formally, given a user query $Q$, a meta-agent $\mathcal{P}$ needs to select some appropriate agents from a total of $n$ agents, denoted as $\mathcal{A}=\{\mathcal{A}_{1},...,\mathcal{A}_{n}\}$. Each agent is associated with a description $d$ that outlines its capability. We denote the collection of all agent descriptions as $\mathcal{D}=\{d_{1},...,d_{n}\}$. The user query $Q$ is decomposed into $m$ sub-tasks, with each sub-task assigned to a specific agent based on their descriptions:
$\mathcal{P}(Q,\mathcal{D},\mathcal{A})=\{(q_{i},\mathcal{A}_{i}^{\prime})\ |\ i\in[m]\}.$ (1)
After that, each selected agent $i\in[m]$ produces a response to its assigned sub-task, denoted as $r_{i}=\mathcal{A}^{\prime}_{i}(q_{i})$. These responses are then utilized to generate the final answer based on $R=\{r_{1},...,r_{m}\}$ to resolve the original query $Q$. Such a process is called agent-oriented planning.
Challenges
[…]
Secondly, assigning sub-tasks to appropriate agents is non-trivial, as the meta-agent can only rely on the agents’ descriptions to determine task allocation in most cases (Shen et al., 2024). However, a concise and highly generalized description that adequately illustrates an agent’s capabilities may not always be available, leading to suboptimal allocation. For example, if the description of a commonsense agent does not specify the extent of its knowledge base, the meta-agent may struggle to ascertain whether it is suitable to assign the task of querying the melting points of tin and copper to that agent, as shown at the lower left of Figure 1.
Figure 1: Examples of agent-oriented planning in multi-agent systems, regarding two challenges (left side) and three design principles (right side).
Design Principles
We further identify three critical principles that guide the design of our framework for effective and efficient agent-oriented planning in multi-agent systems:
  • Solvability. Each sub-task $q_{i}\ \forall i\in[m]$ should be independently and completely resolvable by at least one single agent within the multi-agent system, ensuring that the response for each sub-task can be reliable. If a sub-task does not satisfy solvability, the meta-agent is expected to take some modifications or further decomposition.
  • Completeness. The array of sub-tasks $\{q_{1},...,q_{m}\}$ should include all necessary information from the original user query $Q$, which ensures that the aggregation of responses of these sub-tasks can effectively yield a comprehensive answer to the user query. If the decomposed sub-tasks fail to satisfy completeness, the decomposition process should be revisited by the meta-agent.
  • Non-Redundancy. The array of sub-tasks $\{q_{1},...,q_{m}\}$ should not include redundant elements, avoiding those task executions that are either irrelevant to resolving $Q$, or duplicated. The principle of non-redundancy promotes that the sub-tasks form a minimal effective set necessary to address the user query, enhancing overall efficiency.
Several examples are provided on the right side in Figure 1 for a better understanding of these design principles. In Section 3, we provide more details of the mechanisms in the proposed framework that are designed to fulfill the aforementioned principles.
3 Agent-Oriented Planning in Multi-Agent Systems
In this section, we introduce the details of the proposed agent-oriented planning framework. The overall architecture is illustrated in Figure 2.
Figure 2: Overall architecture of the proposed agent-oriented planning framework.
3.1 Fast Decomposition and Allocation
First of all, following existing studies (Shen et al., 2024; Chen et al., 2023), we provide detailed instructions to the meta-agent to guide it in performing agent-oriented planning. Different from previous studies, the provided instructions incorporate the following requirements: (i) Integration of user query $Q$ and all the agent descriptions $A$: We include both the user query and the descriptions of agents in the instructions, promoting the meta-agent to fully consider the capabilities of each agent and tailor the sub-tasks to align with these capabilities. (ii) Suggestions for assigned agents: We require the meta-agent to provide suggestions for the agents assigned to each decomposed sub-task. Although we tend to separate task decomposition and assignment into two independent tasks performed sequentially, the experimental observations indicate that combining these two tasks enhances the effectiveness of agent-oriented planning. (iii) Structured decomposition of tasks: The decomposed tasks are required to be structured in a sequential manner, specifying any dependencies that may exist between sub-tasks, which ensures that the execution of tasks follows a correct logical order. The adopted prompt can be found in Appendix A.1.
The aforementioned process of agent-oriented planning, which is called fast decomposition and allocation in this study, heavily relies on the capabilities of the meta-agent and meticulously designed system prompts. While such an end-to-end approach can achieve high efficiency, its success rate, namely fulfilling the three design principles mentioned above and providing reliable answers to the original user query, and its stability, namely the meta-agent can follow instructions and produce formatted responses, might not be satisfactory.
Inspired by recent studies on scaling laws during inference (OpenAI, 2024b), we propose to design mechanisms to guide the meta-agent toward more comprehensive reasoning processes. The results of fast decomposition and allocation should be viewed as intermediate outputs rather than final results, which need to be further evaluated to offer detailed revision suggestions for the meta-agent. More details on this proposed approach are provided in the next section.
3.2 Is A Sub-Task Completely Resolved?
To determine whether a sub-task can be resolved by a single agent, i.e., satisfying the principle of solvability, a straightforward solution is to send the sub-task to every agent in the system. Each agent generates a response to the sub-task, allowing us to select the best answer and assess whether it completely resolves the sub-task. However, this approach is often impractical due to the unaffordable overhead. For a user query decomposed into $m$ sub-tasks, a multi-agent system with $n$ agents would need to execute a total of $m\times n$ agent calls for just a single trial, making it an inefficient strategy in scenarios with a large number of agents and sub-tasks.
To tackle this, we propose a reward model designed to provide efficient evaluations of the solvability of sub-tasks, which aims to predict the quality of the agents’ responses to sub-tasks without necessitating actual agent calls.
Specifically, we first prepare a dataset for training the reward model. For a given user query $Q$, we follow the fast decomposition and allocation process introduced in Section 3.1, requiring the meta-agent to decompose $Q$ into several sub-tasks and select $l$ agents for each sub-task based on the agents’ descriptions, which can be formally given as follows:
$\mathcal{P}(Q,\mathcal{D},\mathcal{A})=\{(q_{i},\mathcal{A}_{i,1},...,\mathcal{A}_{i,l})\ |\ i\in[m]\}.$ (2)
After that, we execute the plan, i.e., sending the sub-tasks to the assigned agents, and obtain all the responses from agents as:
$\{r_{i,j}=\mathcal{A}_{i,j}(q_{i})\ |\ i\in[m],j\in[l]\},$ (3)
where the choice of $l$ can be a trade-off. A large $l$ leads to comprehensive responses from lots of agents for constructing the training dataset, while it might need more computation resources and affect the overall quality of the dataset. In this study, we set $l$ to be half the number of agents in the multi-agent systems.
We utilize a scorer $\mathcal{S}$ that evaluates agents’ responses to sub-tasks, i.e., $\mathcal{S}(q_{i},r_{i,j})=s_{i,j}$, which serves as annotations in the training dataset. The scorer $\mathcal{S}$ provides evaluations from three key aspects: correctness, relevance, and completeness, and can be implemented using powerful LLMs (OpenAI, 2024a) or human annotators. We apply this scoring process across a diverse array of user queries $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and collect the results to form the training dataset, denoted as $\mathcal{T}=\{(q_{k,i},d_{k,i,j},s_{k,i,j})\ |\ k\in[K],i\in[m_{k}],j\in[l]\}$. More details on the construction of the dataset, such as the adopted prompts, are summarized in Appendix C.
The reward model, parameterized as $\theta$, consists of embedding layers followed by fully connected layers. We obtain embeddings for both the sub-task and the agent’s description separately, then concatenate these embeddings together and feed them into the fully connected layers for further processing. To provide evaluations without making actual agent calls, we design a training objective aimed at minimizing the discrepancies between the model’s predictions and the annotations provided by the scorer, which can be formulated as follows:
$L(\mathcal{T},\theta)=\frac{1}{K}\sum_{k=1}^{K}\frac{1}{m_{k}}\sum_{i=1}^{m_{k}}\frac{1}{l}\sum_{j=1}^{l}(s_{k,i,j}-\mathcal{M}_{\theta}(q_{k,i},d_{k,i,j}))^{2}.$ (4)
With the well-trained reward model, the evaluation of the results produced by fast decomposition and allocation can be summarized as follows. For each suggestion provided by the meta-agent, stating the assignment of sub-task $q_{i}$ to agent $\mathcal{A}_{i}^{\prime}$, the reward model predicts a score denoted as $\mathcal{M}_{\theta}(q_{i},d_{i}^{\prime})$. A sufficiently high predicted score, which implies that $q_{i}$ can be completely resolved by agent $\mathcal{A}_{i}^{\prime}$, leads to the acceptance of this suggestion. Conversely, if the predicted score falls below a predefined threshold, which indicates that the assignment should be revisited by the meta-agent, the reward model provides a set of scores $\hat{s}_{i,j}=\mathcal{M}_{\theta}(q_{i},d_{j})$ for all the agents $\mathcal{A}_{j},j=1,...,n$. From these scores, the meta-agent reports the optimal solution as $j_{max}=\mathrm{argmax}_{j}\,\hat{s}_{i,j}$.
Note that there may be scenarios where the highest score $\hat{s}_{j_{max}}$ is still below the threshold, indicating that the sub-task $q_{i}$ does not satisfy the principle of solvability and that no single agent can resolve it independently. In these cases, the meta-agent $\mathcal{P}$ is required to perform a replan on $q_{i}$. The prompt used for replan is detailed in Appendix A.2.
3.3 Sub-Task Modifications
In addition to the two scenarios mentioned in Section 3.2, where the assignment of sub-task $q_{i}$ to agent $\mathcal{A}_{i}^{\prime}$ is either predicted to be acceptable by the reward model or to be entirely inappropriate and requires a replan process, there also exists another situation where agent $\mathcal{A}_{i}^{\prime}$ might not provide a reliable response to sub-task $q_{i}$, even though the assignment seems reasonable according to the reward model.
We identify two intrinsic reasons for such a situation in agent-oriented planning. First, the description of the sub-task may lack critical information necessary for solving the problem, such as the interpretation of contextually elements like pronouns or other ambiguous terms. In such cases, while the agent possesses the capability to resolve the sub-task, the missing information can hinder it to provide a satisfactory response. Second, the sub-task $q_{i}$ might be too complex, therefore a single agent can only address parts of this sub-task based on its expertise and tools. In this subsection, we design mechanisms to distinguish between these two cases and propose corresponding solutions.
To be specific, for each agent, we propose to construct a set of representative works that record the tasks it has completely resolved. We would initially bypass the construction and update for these representative works (refer to Section 3.5 for more details) and focus on how to utilize them to make sub-tasks modifications here. Given a sub-task $q_{i}$, we calculate its similarity with a representative work $q_{t}$ following $cos(\mathcal{E}(q_{i}),\mathcal{E}(q_{t}))$, where $\mathcal{E}$ is the embedding part in $\mathcal{M}$ and $cos(u,v)=\frac{u\cdot v}{\left\|u\right\|_{2}\left\|v\right\|_{2}}$ calculates the cosine similarity. We define the similarity as follows:
$sim(q_{i},\mathcal{Q}_{j})=\max\{cos(\mathcal{E}(q_{i}),\mathcal{E}(q_{t}))|q_{t}\in\mathcal{Q}_{j}\},$ (5)
where it can be denoted as $sim_{i,j}$ for short.
A large $\max\{sim_{i,j}\ |\ j\in[n]\}$ indicates that the representative works of the agent corresponding to $\max\{sim_{i,j}\ |\ j\in[n]\}$ contain a sub-task similar to $q_{i}$. Such a case should be attributed to the first reason mentioned above, and we tend to request the meta-agent to perform re-describe on $q_{i}$ according to the similar representative work.
On the other hand, if the calculated similarity does not meet the threshold, we regarded $q_{i}$ as too complex for any single agent in the system to solve (corresponding to the second reason above). As a result, the meta-agent $\mathcal{P}$ is requested to perform plan-in-detail for further decomposing the sub-task into simple ones. To avoid creating redundant sub-tasks, we provide $\mathcal{P}$ with all the sub-tasks and explicit instruction on avoiding overlapping sub-asks. The prompt used for sub-task modification can be found in Appendix A.3 and Appendix A.4.
3.4 Detector for Completeness and Non-Redundancy
Following the idea in fast decomposition and allocation, we attempt to enhance the proposed framework by incorporating detailed instructions to satisfy the principles of completeness and non-redundancy, such as “make sure all the important information of the original task such as nouns or numbers are included in the sub-tasks”. However, empirical observations reveal that these instructions do not mitigate these issues. In fact, more than 15% of user queries still exhibit such issues during decomposition.
To tackle these issues, we involve a detector, implemented with providing a role-play prompt to LLMs, to evaluate both the completeness and non-redundancy of the intermediate results provided by the meta-agent. For the principle of completeness, the detector extracts all key elements and requirements from the original task, and then matches these elements against each sub-task, determining whether the sub-tasks collectively address all essential aspects of the original task. Regarding the principle of non-redundancy, the detector checks for identical information and requirements across sub-tasks. The detector flags the redundancy if two sub-tasks include overlapping content.
When the provided results do not satisfy the principles of completeness or non-redundancy, the detector identifies the missing key information or redundant content, and offers recommendations for refining the plan, such as suggestions for supplementing missing details or removing overlapping sub-tasks.
The prompt used for the detector is provided in Appendix A.5. From the experimental results shown in Section 2, we observe that the detector effectively mitigates the issues of task decomposition incompleteness and redundancy, leading to an effective and efficient agent-oriented planning process.
3.5 Feedback Loop
An automatic feedback loop is integrated into the proposed agent-oriented planning framework, promoting ongoing enhancement of the meta-agent.
As introduced in Section 3.2, the reward model can identify certain sub-tasks as being resolvable by specific agents. These identified sub-tasks are then collected to form the representative works of the corresponding agents. The training data for the reward model, which includes the sub-tasks and the ground truth scores of responses provided by the agents, is utilized to initialize these representative works.
Note that the representative works are continuously updated. When a user query is completely resolved, the sub-tasks decomposed from that user query can be added to the representative works of the agents who are selected to provide response to these sub-tasks. To prevent redundancy in the representative works, we implement a similarity threshold. This threshold ensures that only new and sufficiently distinct sub-tasks are incorporated into an agent’s representative works, maintaining the diversity and relevance of the tasks that each agent has previously resolved.
4 Experiments
In this section, we provide empirical comparisons between the proposed agent-oriented planning framework and the existing studies.
4.1 Settings
Datasets & Evaluations
We conduct experiments based on a numerical reasoning dataset (Kim et al., 2024), which necessitates the collaboration of multiple agents in resolving the queries. For example, a query can be “If Sarah wants to buy one BMW X5 and one Tesla Model 3, how much more would she need to pay to buy the BMW X5 compared to the Tesla Model 3?” To resolve this query, we first need to search for the prices of the BMW X5 and Tesla Model 3, and then calculate the price difference between them. Following previous study (Kim et al., 2024), we construct a training dataset consisting of 1,440 queries, along with an annotated test dataset consisting of 292 queries. Note that the topics in the training dataset and test dataset may vary slightly.
For quantification comparisons, we provide instructions to GPT-4o (OpenAI, 2024b) for assisting in judging whether the execution results align with the ground truth, and calculate the accuracy as the evaluation metrics. The prompt used for evaluation can be found in Appendix D.
Involving Agents
In the experiments, the multi-agent system includes a meta-agent and several diverse agents, including a code agent, math agent, search agent, and commonsense agent. The descriptions of these agents, which are similar to those in previous studies (Kim et al., 2024) for a fair comparison, are shown below:
  • Code Agent: Generate code in Python for precise computations to solve the given task.
  • Math Agent: Answer math questions by reasoning step-by-step.
  • Search Agent: Call the Bing Search API to obtain information related to the given task.
  • Commonsense Agent: Answer the given question using commonsense reasoning.
The meta-agent is tasked with decomposing user queries into sub-tasks and assigning the most suitable agents to execute those sub-tasks. We use GPT-4o as the employed LLM for all agents, and the used prompt are provided in Appendix B.
Reward Model
We adopt all-MiniLM-L6-v2 (Wang et al., 2020) as the embedding layers of the reward model, which maps text sentences or paragraphs into a 384-dimensional dense vector space. The embeddings for both the sub-task and the agent’s description are computed separately, and are concatenate together to form a 768-dimensional dense vector. Following the embedding model, we add a three-layer multilayer perceptron (MLP) with output dimensions of 256,64, and 1, respectively. We freeze the embedding layers and only finetune the parameters of the MLP, making the training process efficiently. The batch size are set to 32, the learning rate is 1e-3. We train the reward model for 50 epochs on one Tesla V100-SXM2-32GB GPU.
Baselines
For the single-agent systems, we utilize GPT-4o as a baseline method, providing user queries directly without any additional prompt engineering. Besides, we provide GPT-4o with Chain-of-Thought (CoT) (Wei et al., 2022) prompt and Zero-Shot CoT (Kojima et al., 2022) prompt to guide LLMs to perform reasoning processes, resulting in two baselines denoted as CoT and Zero-Shot CoT. Taking a further step, we instruct GPT-4o to serve as a meta-agent responsible for performing fast task decomposition and allocations. We implement two baselines, denoted as Meta-Agent and Meta-Agent: Traversal, respectively. Meta-Agent denotes that the LLM selects one agent for each sub-task, while Meta-Agent: Traversal denotes that GPT-4o iteratively queries all agents for each sub-task and selects the best response.
Regarding multi-agent systems, we compare the proposed framework with REACT (Yao et al., 2022) and HUSKY (Kim et al., 2024), which are representative methods involving task decomposition and sub-task execution.
Table 1: Comparisons between the proposed agent-oriented planning framework and baselines.
Method Accuracy (%) Prompt Tokens (M) Completion Tokens (M) Time (s)
GPT-4o 33.3 0.02 0.09 1,968
CoT 35.6 0.07 0.08 1,375
Zero-Shot CoT 32.2 0.02 0.11 1,565
Meta-Agent 30.0 0.65 0.13 5,364
Meta-Agent: Traversal 35.2 3.07 0.69 23,175
REACT 37.6 2.47 0.19 11,510
HUSKY 39.6 0.83 0.15 10,394
Ours 43.7 1.12 0.38 11,869
4.2 Comparisons and Analysis
The comparisons between the proposed agent-oriented planning framework and the baseline methods are shown in Table 1. We adopt accuracy as the metric for comparing effectiveness, and use prompt tokens, completion token, and the execution time for comparing efficiency.
Overall, the experimental results demonstrate that the proposed agent-oriented planning framework achieves notable improvements compared to all baseline methods. To be specific, when comparing with single-agent systems including GPT-4o, CoT, and Zero-Shot CoT, the proposed agent-oriented planning framework outperforms these systems by 10.4%, 8.1%, and 11.5% in terms of accuracy, respectively. These improvements can be attributed to the collaboration among multiple diverse agents and the effective scheduling provided by the meta-agent. It is not surprising to observe that the proposed framework incurs higher resource costs and longer inference times compared to single-agent systems. These additional costs during the inference phase, which includes task decomposition, allocation, and modifications carried out by the meta-agent, are affordable and can be worthwhile as long as they bring significant improvements in accuracy and stability when applying real-world applications.
Compared to systems that also involve a meta-agent for task decomposition and allocation, we can observe from the table that the proposed method achieves at least a 4.1% improvement in terms of accuracy while maintaining the same level of computation costs and inference times. The results of Meta-Agent and Meta-Agent: Iteration indicate that simply instructing GPT-4o to perform task decomposition and allocation does not always yield satisfactory responses, even though the capabilities of LLM are recognized as powerful. Existing studies in task composition and allocation fail to consider the abilities and characteristics of the agents beyond their descriptions, and lack mechanisms for modifications and feedback within the multi-agent system, which can be hindered by the challenges as summarized in Section 2. The proposed framework is built upon these identified design principles and incorporates novel mechanisms for automatically evaluating intermediate results, allowing timely modifications and revisions, which guides and leverages the strengths of LLMs within the multi-agent system for effectively resolving user queries and enhancing overall system performance.
4.3 Further Discussions
Ablation Study
We conducted an ablation study to confirm the contributions of different components in the proposed framework. Specifically, we disable the detector, the reward model, and the representatives of agents in separate experiments, with results reported in Table 2. From these results, we can observe that the detector has a significant impact on execution accuracy, as indicated by a notable increase in the incompleteness rate during the meta-agent’s task decomposition phase. Both the reward model and the representative works are necessary in ensuring the solvability of tasks and in selecting the most suitable agents for each sub-task. Overall, these results demonstrate that all three components are indispensable, working together to ensure the feasibility of task decomposition and allocation, leading to satisfactory responses to the original user queries.
Table 2: Experimental results of ablation study.
Method Accuracy (%) Prompt Tokens (M) Completion Tokens (M) Time (s)
Ours 43.7 1.12 0.38 11,869
w/o Plan Detector 36.6 1.05 0.29 9,504
w/o Reward Model 38.7 1.17 0.37 11,651
w/o Representative Works 41.1 1.14 0.36 11,178
Impact of Scorer and Reward Model
The training dataset is constructed based on annotations provided by the scorer, which in the previous experiments is implemented using LLMs. In this section, we shift to using human annotators to manually provide scores for the responses, evaluating whether a reward model trained on this manually labeled dataset would further enhance performance. This approach is denoted as Manual Scoring. More detailed information about the manual scoring process can be found in Appendix C. Besides, based on the manually scored training dataset, we investigate the setting of updating the embedding layers of the reward model, denoted as Full Parameter Tuning.
The experimental results shown in Table 3 indicate that both manual scoring and full parameter tuning of the reward model lead to improved performance. These results suggest the potential for further improving the proposed agent-oriented planning framework by providing high-quality datasets for training robust reward models.
Table 3: Experimental results on the impact of scorer and reward model
Method Accuracy (%) Prompt Tokens (M) Completion Tokens (M) Time (s)
Ours 43.7 1.12 0.38 11,869
+ Manual Scoring 46.9 1.26 0.40 12,676
+ Full Parameter Tuning 47.9 1.28 0.41 13,063
5 Related Work
The field of artificial general intelligence has witnessed a transformative development with the advent of large language models (LLMs), such as GPT-3 (Brown, 2020), GPT-4 (Achiam et al., 2023), PaLM (Chowdhery et al., 2023), LLaMa (Touvron et al., 2023) and so on (Devlin et al., 2018; Zhang et al., 2022; Chowdhery et al., 2023; Le Scao et al., 2023). LLMs have demonstrated remarkable capabilities in understanding and generating human language (Zhao et al., 2023), and solving tasks such as mathematical problems and commonsense reasoning (Imani et al., 2023; Ahn et al., 2024). To further enhance the ability of LLMs in handling complex problems, some research focuses on improving the internal reasoning capabilities of LLMs, typically involves decomposing complex questions into sequential intermediate steps before generating the final answer, as exemplified by Chain-of-Thought (CoT) prompting (Wei et al., 2022) and its variants (Kojima et al., 2022; Wang et al., 2022; Zhou et al., 2022). Some methods opt to generate a dozen possible paths to comprise the candidate plan set (Yao et al., 2024; Besta et al., 2024; Zhao et al., 2024; Hao et al., 2023) and select the optimal plan with strategies like heuristic search algorithms (Xiao & Wang, 2023). In environments with complex constraints, such as generating permissible actions, several methods combine LLMs with external planners such as symbolic planners (Liu et al., 2023; Dagan et al., 2023; Guan et al., 2023) and neural planners (Yao et al., 2020; Lin et al., 2024).
Despite these advancements, LLMs still face challenges in addressing tasks that require extensive knowledge, such as avoiding hallucinations, integrating diverse domain knowledge, and executing long-horizon planning (Qin et al., 2023a; Bubeck et al., 2023; Maynez et al., 2020). To address these issues, recent work has focused on multi-agent systems, which leverage the collective intelligence and specialized skills of multiple LLM-empowered agents to solve tasks. Compared with reinforcement learning, LLM-empowered agents have more comprehensive internal world knowledge, and can provide natural language interfaces to interact with humans, which is more flexible and explainable (Wang et al., 2024). Serving as the main controller or “brain” that manages the flow of operations, the meta-agent has gained the most significant attention. The meta-agent is responsible for understanding the overall task, decomposing it into smaller ones, and coordinating the sub-tasks to the agents. There are mainly two technical approaches in task decomposition: interleaved decomposition and decomposition-first. In the first kind of method, React (Yao et al., 2022) generates reasoning and action in an interleaved manner, where reasoning helps the meta-agent update action plans and action enables interaction with agents. Visual ChatGPT (Wu et al., 2023a) meticulously designs a series of prompts to help inject the visual information into ChatGPT, designed to solve complex visual questions with diverse visual foundation models step-by-step. HUSKY (Kim et al., 2024) trains an action model to iterate between two stages: generating the next action to take and executing the action using expert models and updating the current solution state.
Regarding complex tasks, excessively long trajectories might bring hallucinations. Given that, HuggingGPT (Shen et al., 2024) carefully designs prompts to instruct ChatGPT to decompose tasks, selecting models in HuggingFace and summarizing the response. AutoAgents (Chen et al., 2023) dynamically generates multiple required agents based on task content and plans solutions for the current task based on the generated expert agents. Chameleon (Lu et al., 2024) prompts the meta-agent with tool descriptions and usage examples to infer a program composed of a sequence of tools to execute in order to generate the final response.
Although remarkable progress has been made, plans generated solely through prompt-based methods often fall short of meeting the principles of solvability, completeness, and non-redundancy—issues that have largely been overlooked in prior work. Besides, the meta-agent’s ability to make suitable schedules can be hindered when relying only on agents’ descriptions without considering their actual capabilities or past performance. To address these issues, we propose a novel framework to perform task decomposition and allocation satisfying the aforementioned three principles.
6 Conclusion
In this study, we propose a novel agent-oriented planning framework for multi-agent systems, following three critical design principles to ensure that the meta-agent can effectively decompose the user query into several sub-tasks for producing satisfactory responses. The proposed framework utilizes a fast decomposition and allocation process, which relies on the ability of LLMs to generate an intermediate schedule efficiently. After that, a reward model and the representative work mechanism are employed to evaluate these intermediate results, routing three different paths for executing the sub-task or making necessary modifications to align the sub-task with agents, such as replan the sub-task, plan in detail, and re-describe. Extensive experiments demonstrate that the proposed agent-oriented framework achieves significant improvements over both the existing single-agent and multi-agent baseline methods. We provide discussions on the contributions of different components in the framework and the potential for improvements in agent-oriented planning. We will release the source code to promote further research in the community.
Appendix A Plan Generation and Modification
A.1 Meta-agent Prompt
The one-shot prompt for meta-agent’s fast decomposition and allocation is shown in Figure 3.
meta_agent_prompt = ’’’
You are a planning agent. You are responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task. Your main goal is to efficiently and accurately complete task planning based on the descriptions of agents provided, ensuring the coherence and quality of the sub-tasks.
Please output the sub-tasks and corresponding agents in the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agents; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
The available agents and the corresponding descriptions are: [code_agent: Generate code in Python for precise computations to solve the given task. math_agent: Answer math questions by reasoning step-by-step. search_agent: Call Bing Search API for obtaining information regarding the given task. commonsense_agent: Answer the given question using commonsense reasoning.].
Here is an example:
[…]
Output:
"task": "Determine the population of China.",
"id": 1,
"name": "search_agent",
"reason": "The search_agent can find the most recent and accurate population data for China.",
"dep": []
"task": "Calculate 1% of the population of China.",
"id": 2,
"name": "math_agent",
"reason": "The math_agent can reason through the calculation step-by-step.",
"dep": [1]
"task": "Determine the number of full flights needed to transport 1% of the population of China to Indonesia, given that each plane can carry 300 passengers.",
"id": 3,
"name": "math_agent",
"reason": "The math_agent can reason through the division and rounding process.",
"dep": [2]
Given the user query, output the task plan with the format above directly. Make sure all the important information such as nouns or numbers are included in the sub-tasks. If you think the query can be done with just one agent, you can output only one sub-task.
Figure 3: One-shot prompt for the meta-agent.
A.2 Replan
The prompt for replanning is shown in Figure 4.
replan_prompt = ’’’
You are a planning agent. You are preliminarily responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task according to the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agents; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
The available agents and the corresponding descriptions are: [code_agent: Generate code in Python for precise computations to solve the given task. math_agent: Answer math questions by reasoning step-by-step. search_agent: Call Bing Search API for obtaining information regarding the given task. commonsense_agent: Answer the given question using commonsense reasoning.].
Given the user query: %s, the preliminary task decomposition is: %s.
But the sub-task: %s cannot be solved by any agent. Now you are responsible for replaning this sub-task based on agents capabilities. Output the new sub-task with the format above directly.
Figure 4: Prompt for replanning.
A.3 Plan in Detail
The prompt for planning in detail is shown in Figure 5.
plan_in_detail_prompt = ’’’
You are a planning agent. You are preliminarily responsible for decomposing the given query into sub-tasks and choose the most suitable agent for each sub-task according to the following json format: [{"task": task_description, "id": task_id, "name": name_of_agent, "reason": your_detailed_reason_for_the_choice, "dep":dependency_task_ids}]. In this format, "task" is the description of the sub-task, which will be used as the input of the chosen agents; "dep" denotes the id of the previous sub-task which generates a new resource relied by the current sub-task.
The available agents and the corresponding descriptions are: [code_agent: Generate code in Python for precise computations to solve the given task. math_agent: Answer math questions by reasoning step-by-step. search_agent: Call Bing Search API for obtaining information regarding the given task. commonsense_agent: Answer the given question using commonsense reasoning.].
Given the user query: %s, the preliminary task decomposition is: %s.
But the sub-task: %s cannot be solved only with agent %s. Now you are responsible for planning this sub-task in detail and choose the most suitable agents based on agents capabilities. Output the new sub-tasks with the format above directly. Make sure that there are no duplicate content between new sub-tasks and the given preliminary task decomposition.
Figure 5: Prompt for planning in detail.
A.4 Re-describe
The prompt for re-describing the sub-task is shown in Figure 6.
redescribe_subtask_prompt = ’’’
Rewrite the following sentence based on the given example, while keeping the key information unchanged. Besides, output the rewritten sentence in the form like ***rewritten***.
Here is an example:
Example sentence: Determine the population of the United States in 2022.’
Sentence to be rewritten: Assess the population of China in 2022.’
Rewritten sentence: Determine the population of China in 2022.’
Output: ***’Determine the population of China in 2022.’***
Example sentence: %s
Sentence to be rewritten: %s
Figure 6: Prompt for re-describing the sub-task.
A.5 Detector
The prompt for the detector is shown in Figure 7.
plan_detector_prompt = ’’’
You are a plan detector responsible for analyzing the completeness and redundancy of the plan. Given the query and the plan formulated to solve the query, which involves several sub-tasks, you should do the following things:
1. **Detect whether the plan satisfies the completeness.**: Evaluate whether the set of subtasks covers all key aspects of the original task including important numbers and nouns. Specifically, check if each important element and requirement from the original task is addressed by at least one subtask. Provide a brief explanation if any key information is missing.
2. **Detect whether the plan satisfies the non-redundancy.**: Evaluate whether any two sub-tasks contain identical information and requirements. If there is any redundant part, list and provide suggestions for optimizing the plan.
For example:
Task: If a plane can carry 300 passengers and flies from Brazil to Nigeria with a full load, then returns with only 75% capacity filled, how many passengers in total has it transported between the two countries in one round trip?
Subtask 1: Determine the number of passengers transported from Brazil to Nigeria in one flight with a full load. Dependency: []
Subtask 2: Determine the number of passengers transported from Nigeria to Brazil in one flight with 75% capacity filled. Dependency: []
Subtask 3: Calculate the total number of passengers transported between Brazil and Nigeria in one round trip. Dependency: [1, 2]
Analyse: This plan does not satisfy completeness because the subtask loses the information of a plane can carry 300 passengers of the original task. This plan satisfies non-redundancy because each subtask has a unique focus and there is no overlap in the information covered.
Suggestions: Add the information of a plane can carry 300 passengers to subtask 1 and subtask 2.
If there is no need to modify the plan, just return The plan satisfies completeness and non-redundancy.’.
Figure 7: Prompt for the detector.
Appendix B Agents
The prompts for the code, math, search and commonsense agent are shown in Figure 8, 9, 10, 11 separately.
code_agent_prompt = ’’’You ara a code agent. You can be used for : 1) computing large numbers, fractions or decimals. 2) counting or averaging long lists of numbers. 3) performing date-related operations, such as counting the number of days between two dates. Write code in Python to solve the given task with history. Give the code in the following form directly.
- Here is an example:
Task: Calculate the combined population of China and India in 2022.
History: The answer of Determine the population of China in 2022’ is 1.412B. The answer of Determine the population of India in 2022’ is 1.417B.
Code:
‘‘‘python
# Given populations
population_china_2022 = 1.412 * 10**9 # 1.412 billion
population_india_2022 = 1.417 * 10**9 # 1.417 billion
# Calculate combined population
combined_population_2022 = population_china_2022 + population_india_2022
# Print the result
print(f"The combined population of China and India in 2022 is {combined_population_2022} people.")
Task: %s
History: %s
Code:
rewrite_code_agent_prompt = ’’’You are a rewrite agent. Given the input question, the code addressing this question and the corresponding output, rewrite the output into a complete sentence that integrates information from the question and the code output.
Question: %s
Code: %s
Code output: %s
Output:
Figure 8: Prompt for the code agent.
[…]
Question: %s
History: %s
Solution:
Figure 9: Prompt for the math agent.
search_agent_prompt = ’’’You are a search agent. Write a concise, informative Bing Search query for obtaining information regarding the given task.
- Here is an example:
Task: Determine the population of China in 2022.
History: None
Search query: China population 2022
Task: %s
History: %s
Search query:
rewrite_search_agent_prompt = ’’’You are a rewrite agent. Given the search question, the response in the web_pages from the bing search api, answer the search question with the information from the response in concise words. Remove redundant information that is irrelevant to the question.
Question: %s
Answer_box: %s
Answer:
Figure 10: Prompt for the search agent.
commonsense_agent_prompt = ’’’You are a commonsense agent. You can answer the given question with logical reasoning, basic math and commonsense knowledge.
Question: %s
History: %s
Solution:
Figure 11: Prompt for the commonsense agent.
Appendix C Constructing Training Dataset
The prompt for the scorer is shown in 12. The criteria for manual scoring are consistent with the standards outlined in the prompt provided to the scorer.
scorer_prompt = ’’’
Please act as an impartial judge and evaluate the quality of the response provided by the %s to the user task. Your evaluation should consider three factors: correctness, relevance and completeness. Assign a score of 0, 1 or 2 for each factor and provide a brief explanation for your score. The following is the grading criteria.
Correctness
0: The response contains severe errors and is completely inaccurate.
1: The response has some errors, but the main content is generally correct
2: The response is completely accurate and fully meets the requirements of the task.
Relevance
0: The response is minimally relevant to the task and completely off-topic.
1: The response is somewhat relevant to the task but may include some unrelated content.
2: The response is highly relevant to the task, directly addressing the core issue without any unrelated content or deviation.
3. Completeness
0: The response lacks necessary detail or key information, resulting in an incomplete understanding of the task.
1: While the response addresses part of the task, more information or content is needed for completeness.
2: The response provides comprehensive information and detailed explanations.
Besides, summarize the final result in the form like ’**Correctness: score, Relevance: score, Completeness: score**’ at the end of your response, where score can be chosen from 0, 1 and 2.
Task
%s
[The Start of Agents Response]
%s
[The End of Agents Response]
Figure 12: Prompt for the scorer.
We design a mapping function that converts different combinations of correctness, relevance, and completeness values into a final score, which is shown in Figure 13. Then, a score of 7 or higher is considered sufficiently high, while a score of 1 or lower is considered sufficiently low.
score_map = {
(2, 2, 2): 8,
(2, 1, 2): 7,
(2, 2, 1): 6,
(2, 1, 1): 5,
(1, 2, 2): 4,
(1, 1, 2): 3,
(1, 2, 1): 2,
(1, 1, 1): 1,
def level_score(correctness, relevance, completeness):
return score_map.get((correctness, relevance, completeness), 0)
Figure 13: Mapping function.
Appendix D Evaluation
The prompt for comparing the response with the ground truth is shown in Figure 14.
evaluate_prompt = ’’’
You are CompareGPT, a machine to verify the correctness of predictions. Answer with only yes/no.
You are given a question, the corresponding ground-truth answer and a prediction from a model. Compare the "Ground-truth answer" and the "Prediction" to determine whether the prediction correctly answers the question. The prediction may contain extra information, but a correct prediction includes the ground-truth answer. You can answer "yes" if the prediction includes the ground-truth answer. You must answer "no" if there are any specific details in the ground-truth answer that are not mentioned in the prediction. If the prediction states something as a possibility, treat it as a definitive answer. Note that the error within three decimal places is negligible.
Question: %s
Ground-truth answer: %s
[Start of the prediction]
%s
[End of the prediction]
Figure 14: Prompt for comparing the response with the ground truth.